OpenStudy (anonymous):

Given that 3sin^2-8cosx-7=0 show that, for real values of x cos x = -2/3 I'm really bad at trig identities so if anyone could help it would be great appreciated :D

hartnn (hartnn):

did u know that \(\sin^2x+\cos^2x=1\) ?

OpenStudy (anonymous):

Yes I do

hartnn (hartnn):

so, from that what does sin^2 x = ... ?

OpenStudy (anonymous):

1-cos^2x

hartnn (hartnn):

good! so, plug in 1-cos^2 x in place of sin^2 x in 3sin^2 x-8cos x -7=0 what u get ?

OpenStudy (anonymous):

3(1-cos^2)-8cosx-7=0?

hartnn (hartnn):

simplify that a bit ? to bring it in the form of \(ax^2+bx+c=0\) ?

OpenStudy (anonymous):

3-3cos^2-8cosx-7=0 => -3cos^2-8cos-4=0

hartnn (hartnn):

cool, lets put cos x = X for simplicity can you solve this quadratic ? -3X^2 -8X-4 =0 ??? or 3X^2+8X+4=0 ?

OpenStudy (anonymous):

(-3x-2)(x-2)=0?

hartnn (hartnn):

absolutely correct! good work :)

hartnn (hartnn):

so, x was cos x -3cos x-2 = 0 gives you what ?

OpenStudy (anonymous):

-3cosx=2 => cosx=-2/3 Thanks so much you make the questions so much easier for me :)

hartnn (hartnn):

cos x+2 =0 gives cos x = -2 but the range of cos is between 1 and -1 so for no value of x. cos x will be -2

hartnn (hartnn):

and you are welcome ^_^

OpenStudy (anonymous):

Thanks!